OpenStudy (anonymous):

is f(x) = 3root2x-1 a rational function, as well as a polynomial function? (I know it's radical.

OpenStudy (anonymous):

oh it's one to one as well right?

OpenStudy (anonymous):

oh and how the heck to do I solve for it's inverse 0.0 s

jimthompson5910 (jim_thompson5910):

I'm assuming that it's \[f(x)=3\sqrt{2x-1}\] This is a radical function (since it has a radical in it), but it is NOT rational (since it's not a ratio of two polynomials). This function is one-to-one and you can show that by proving that f(x)=f(y) <--> x = y Finding the inverse: \[\large f(x)=3\sqrt{2x-1}\] \[\large y=3\sqrt{2x-1}\] \[\large x=3\sqrt{2y-1}\] \[\large \frac{x}{3}=\sqrt{2y-1}\] \[\large \frac{x^2}{9}=2y-1\] \[\large \frac{x^2}{9}+1=2y\] \[\large \frac{x^2+9}{9}=2y\] \[\large \frac{x^2+9}{18}=y\] \[\large y=\frac{x^2+9}{18}\] So the inverse is \[\large f^{-1}(x)=\frac{x^2+9}{18}\]